Complete the following functions
F(x)=x-2, g(x)=x^2
Find (f+g)(x) and its domain
find fg(x) and its domain
find (F/g)(x) and its domain.
Can someone please show show a step by step instructions for at least one of them please.

Answer by Boreal(15235)

(f+g)x=
x-2 +x^2 +2= x^2 +x domain is all reals.
fg(x)= (x-2) (x^2)=x^3-2x^2 domain is all reals.
(f/g) x= (x-2)/x^2, and that cannot be simplified. Domain is all reals.